### v2

# Exponential Sums by Irrationality Exponent

Abstract.
In this article, we give an asymptotic bound for the exponential sum of the Möbius function $\sum_{n\leq x}\mu(n)e(\alpha n)$ for a fixed irrational number $\alpha\in\mathbb{R}$. This exponential sum was originally studied by Davenport and he obtained an asymptotic bound of $x(\log x)^{-A}$ for any $A\geq 0$. Our bound depends on the irrationality exponent $\eta$ of $\alpha$. If $\eta\leq 5/2$, we obtain a bound of $x^{4/5+\varepsilon}$ and, when $\eta\geq 5/2$, our bound is $x^{(2\eta-1)/2\eta+\varepsilon}$. This result extends a result of Murty and Sankaranarayanan, who obtained the same bound in the case $\eta=2$.
Key words and phrases:
exponential sum, Vinogradov bound, irrationality exponent
2020 Mathematics Subject Classification:
Primary 11L07; Secondary 37E10
B. Cha was supported by the Hongik University new faculty research fund. D. H. Kim was supported by National Research Foundation of Korea (RS-2023-00245719).
1. Introduction
Let $e(\alpha)=\exp(2\pi\alpha\sqrt{-1})$. Using Vinogradov’s earlier work [V], Davenport [D] established the following bound for the exponential sum of the Möbius function: for a fixed irrational $\alpha$,
(1) $\sum_{n\leq x}\mu(n)e(\alpha n)\ll_{A}x(\log x)^{-A},$
for any $A\geq 0$, where the implied constant depends only on $A$.
It is possible to improve the above bound at least conditionally, by introducing an assumption on a zero-free region of Dirichlet $L$-functions. Hajela and Smith [HS87] showed that
(2) $\sum_{n\leq x}\mu(n)e(\alpha n)\ll_{\varepsilon}x^{(a+2)/3+\varepsilon},$
under the assumption that
(3) $L(s,\chi)\text{ has no zeros in the half plane }\Re(s)>a.$
Improving (2), Baker and Harman [BH] proved, under the assumption (3) for all Dirichlet $L$-functions, that
(4) $\sum_{n\leq x}\mu(n)e(\alpha n)\ll_{\varepsilon}x^{b+\varepsilon},$
where
$b=\begin{cases}a+\tfrac{1}{4}&\tfrac{1}{2}\leq a<\tfrac{11}{20},\\ \tfrac{4}{5}&\tfrac{11}{20}\leq a<\tfrac{3}{5},\\ \tfrac{a+1}{2}&\tfrac{3}{5}\leq a<1.\end{cases}$
Note that, if we assume GRH for all Dirichlet $L$-functions, the bound of Hajela and Smith becomes $x^{\frac{5}{6}+\varepsilon}$, while that of Baker and Harman is $x^{\frac{3}{4}+\varepsilon}$. More recently, assuming (3) in the range $a\in[1/2,4/7]$, Zhang [Zha] improved Baker and Harman’s bound in this range by proving (4) where
$b=\frac{8a-7a^{2}}{4-2a}.$
For more results on this topic, see [JL19] and the introduction therein.
In this paper, we take a different approach by considering the irrationality exponent of $\alpha$ without any assumption on a zero-free region of $L$-functions. Recall, from [Bugeaud] for example, that the irrationality exponent of $\alpha$ is defined to be the supremum of the set of real numbers $\tau$ for which the inequality
(5) $0<\left|\alpha-\frac{p}{q}\right|<\frac{1}{q^{\tau}}$
is satisfied by infinitely many integer pairs $(p,q)$ with $q>0$. The irrationality exponent is greater than or equal to 2 for any irrational $\alpha$. Our main theorem is
Theorem 1.
Suppose that $\alpha$ is a fixed irrational number whose irrationality exponent $\eta$ is finite. Let $\varepsilon>0$. If $2\leq\eta\leq\frac{5}{2}$, then
$\sum_{n\leq x}\mu(n)e(\alpha n)\ll_{\varepsilon}x^{\left(\frac{4}{5}+\varepsilon\right)}.$
If $\eta\geq\frac{5}{2}$, then
$\sum_{n\leq x}\mu(n)e(\alpha n)\ll_{\varepsilon}x^{\left(\frac{2\eta-1}{2\eta}+\varepsilon\right)}.$
When $\eta=2$, that is, when $\alpha$ is of irrational type 1, Theorem 1 has been already proved in 2002 by Murty and Sankaranarayanan (see Corollary 4 in [MS02]).
Note that the bounds for the exponential sum given by the other authors we quoted above are all uniform, in the sense that they are independent of $\alpha$. Thus what we achieve in this paper is to improve Davenport’s bound at the expense of making it dependent on the irrationality exponent of $\alpha$.
A bound for the exponential sum for the Möbius function can be used to prove Sarnak’s Möbius disjointness conjecture for the simplest type of dynamical system, namely the rotation of $\mathbb{R}/\mathbb{Z}$ by $\alpha$. To recall the statement of the conjecture [Sar] in general, let $(X,T)$ be a topological dynamical system. For any $x\in X$ and $f\in C(X)$, the conjecture asserts that, if $(X,T)$ is determinisitic, namely, of entropy zero, then
(6) $\sum_{n\leq N}\mu(n)f(T^{n}x)=o(N)$
as $N\to\infty$. If $X=\mathbb{R}/\mathbb{Z}$ and if $T=T_{\alpha}$ is the rotation by $\alpha$, that is, $T_{\alpha}(x):=x+\alpha$, the verification of (6) boils down to showing the orthogonality of $\mu(n)$ against $e(\alpha n)$. In fact, one can easily deduce from (1) that the left-hand side of (6) is the big-$O$ of $N(\log N)^{-A}$ (whose implicit constant depends on $x,f$ and $A$), hence $o(N)$. This will in turn yield a better convergence rate with respect to the conjecture (6) of Sarnak. We note here that the conjecture (6) has been proven for several other kinds of systems $(X,T)$, such as Thue-Morse sequences [MR], nilsequences [GT], and horocycle flows [BSZ].
Our proof begins with a few preliminary lemmas in §2, culminating in Proposition 4. The results in §2 are originally due to Vinogradov [V] and Vaughan [Vau]. Our exposition mostly follows [IK] and [GT2]. The proof of Theorem 1 is deduced from Proposition 4.
2. Lemmas
We begin with the, so-called, Vaughan’s identity [Vau], following the exposition in Chapter 13 of [IK] and Section 4 of [GT2]: for fixed $M$ and $N$,
(7) $\sum_{m\leq x}\mu(m)e(\alpha m)=-T_{\mathrm{I}}+T_{\mathrm{II}}+O(\max\{M,N\}),$
where
$T_{\mathrm{I}}:=\underset{\begin{subarray}{c}\ell k\leq x\\ k\leq MN\end{subarray}}{\sum\sum}e(\alpha\ell k)\sum_{\begin{subarray}{c}n\mid k\\ n\leq N\end{subarray}}\mu(n)\mu(k/n)$
and
$T_{\mathrm{II}}:=\underset{\begin{subarray}{c}kn\leq x\\ k>M,n>N\end{subarray}}{\sum\sum}\mu(n)e(\alpha kn)\sum_{\begin{subarray}{c}m\mid k\\ m>M\end{subarray}}\mu(m).$
To prove (7), write
$\mu(m)=\underset{bc|m}{\sum\sum}\mu(b)\mu(c)$
and break up the summation into four regions: (i) $b\leq M$ and $c\leq N$ (ii) $b\leq M$ and $c>N$ (iii) $b>M$ and $c\leq N$ (iv) $b>M$ and $c>N$. Then it is easy to see that $\sum_{(\mathrm{ii})}=\sum_{(\mathrm{iii})}=-\sum_{(\mathrm{i})}$, if $m>\max\{M,N\}$. Thus,
$\sum_{m\leq x}\mu(m)e(\alpha m)=-\underset{\begin{subarray}{c}\ell mn\leq x\\ m\leq M,n\leq N\end{subarray}}{\sum\sum\sum}\mu(m)\mu(n)e(\alpha\ell mn)\\ +\underset{\begin{subarray}{c}\ell mn\leq x\\ m>M,n>N\end{subarray}}{\sum\sum\sum}\mu(m)\mu(n)e(\alpha\ell mn)+O(\max\{M,N\}).$
From this, a simple change of variable changes the first summation in the right hand side above as $T_{\mathrm{I}}$ and the second as $T_{\mathrm{II}}$, establishing (7).
To bound $T_{\mathrm{I}}$ and $T_{\mathrm{II}}$, we write
$T_{\mathrm{I}}=\underset{\begin{subarray}{c}\ell k\leq x\\ k\leq MN\end{subarray}}{\sum\sum}e(\alpha\ell k)\gamma(k,N),$
and
$T_{\mathrm{II}}=\underset{\begin{subarray}{c}kn\leq x\\ k>M,n>N\end{subarray}}{\sum\sum}\mu(n)e(\alpha kn)\tau(k,M),$
where
$\gamma(k,N):=\sum_{\begin{subarray}{c}n\mid k\\ n\leq N\end{subarray}}\mu(n)\mu(k/n),\qquad\tau(k,M):=\sum_{\begin{subarray}{c}m\mid k\\ m>M\end{subarray}}\mu(m).$
Let $d(k):=\sum_{u|k}1$ be the divisor function, that is, the number of positive divisors of $k$. Then it is well known that
$d(k)\ll_{\varepsilon}k^{\varepsilon}.$
Moreover, we clearly have
$|\gamma(k,N)|\leq d(k),\qquad|\tau(k,N)|\leq d(k).$
Now we appeal to the following two lemmas from [IK]. Suppose that $q$ satisfies
$\left|\alpha-\frac{a}{q}\right|\leq\frac{1}{q^{2}}\text{ with }(a,q)=1.$
Lemma 2 (Lemma 13.7 in [IK]).
$\sum_{1\leq m\leq M}\left|\sum_{1\leq n\leq\frac{x}{m}}e(\alpha mn)\right|\ll\left(M+\frac{x}{q}+q\right)\log(2qx)$
as $x\to\infty$.
This proves that
$T_{\mathrm{I}}\ll_{\varepsilon}\left(MN+\frac{x}{q}+q\right)x^{\varepsilon}\log(2qx).$
For $T_{\mathrm{II}}$, we use
Lemma 3 (Lemma 13.8 in [IK]).
Let $\{\alpha_{n}\}$ and $\{\beta_{m}\}$ be sequences of complex numbers whose absolute values are bounded by 1. Then,
$\underset{\begin{subarray}{c}mn\leq x\\ m>M,n>N\end{subarray}}{\sum\sum}\alpha_{m}\beta_{n}e(\alpha mn)\ll\left(\frac{x}{M}+\frac{x}{N}+\frac{x}{q}+q\right)^{\frac{1}{2}}x^{\frac{1}{2}}(\log x)^{2},$
as $x\to\infty$.
Likewise,
$T_{\mathrm{II}}\ll_{\varepsilon}\left(\frac{x}{M}+\frac{x}{N}+\frac{x}{q}+q\right)^{\frac{1}{2}}x^{\frac{1}{2}+\varepsilon}(\log x)^{2}.$
We summarize what we proved as
Proposition 4.
Let $\varepsilon>0$ be given. Then, for fixed $M$ and $N$
(8) $\sum_{m\leq x}\mu(m)e(\alpha m)=-T_{\mathrm{I}}+T_{\mathrm{II}}+O(\max\{M,N\}),$
where $T_{\mathrm{I}}$ and $T_{\mathrm{II}}$ satisfy
(9) $T_{\mathrm{I}}\ll_{\varepsilon}\left(MN+\frac{x}{q}+q\right)x^{\varepsilon}\log(2qx)$
(10) $T_{\mathrm{II}}\ll_{\varepsilon}\left(\frac{x}{M}+\frac{x}{N}+\frac{x}{q}+q\right)^{\frac{1}{2}}x^{\frac{1}{2}+\varepsilon}(\log x)^{2}.$
3. Proof of Theorem 1
Let $\eta$ be the irrationality exponent of $\alpha$ and fix a real $\tau$ with $\tau>\eta$. Then by the definition of irrationality exponent,
$\frac{1}{q^{\tau}}<\left|\alpha-\frac{p}{q}\right|$
is satisfied for all but finitely many integer pairs of $p$ and $q$. Then there exists $i_{0}=i_{0}(\tau,\alpha)$ such that, for all $i\geq i_{0}$,
$\frac{1}{q_{i-1}^{\tau}}<\left|\alpha-\frac{p_{i-1}}{q_{i-1}}\right|<\frac{1}{q_{i}q_{i-1}},$
where $p_{i}/q_{i}$ is the $i$-th principal convergent in the continued fraction of $\alpha$. This implies
(11) $q_{i}<{q_{i-1}}^{\tau-1}$
for all $i\geq i_{0}$. Now, given large $x$, choose $i$ so that ${q_{i-1}}^{\tau-1}\leq x^{(\tau-1)/\tau}<{q_{i}}^{\tau-1}$. Then $q:=q_{i}$ satisfies
(12) $q^{\frac{\tau}{\tau-1}}<x<q^{\tau}.$
From this, we deduce
(13) $q<x^{1-\frac{1}{\tau}},\qquad\text{and}\qquad\frac{x}{q}<x^{1-\frac{1}{\tau}}.$
We now invoke Proposition 4 with $M=N=x^{\frac{2}{5}}$. From (13) and (9),
(14) $T_{\mathrm{I}}\ll_{\varepsilon}\left(x^{\frac{4}{5}}+x^{1-\frac{1}{\tau}}\right)x^{\varepsilon},$
and also from (10)
(15) $T_{\mathrm{II}}\ll_{\varepsilon}\left(x^{\frac{8}{5}}+x^{2-\frac{1}{\tau}}\right)^{\frac{1}{2}}x^{\varepsilon}.$
Combining (14) and (15), we complete the proof of Theorem 1.
Acknowledgements
The authors are thankful to Peter Humphries who brought [MS02] to the authors’ attention.